1. The density of water is 1.00 g/ML at 4 degrees Celsius. How many water molecules are present in 2.56 mL at this temperature?

2. For a 4.35 g sample of ammonia gas (NH3), Determine:

(a) Number of ammonia molecules.
(b) Total number of atoms in a sample.
(c) Volume of the gas at room temperature of 25 degrees Celsius and a pressure of 1 atm.
1
Expert's answer
2020-08-28T09:33:21-0400

Problem 1:

Density "\rho=\frac{m}{V}"

"\nu=\frac{m}{M}=\frac{N}{N_A}" whence "N=\frac{m}{M}N_A=\frac{\rho*V}{M}*N_A"

Molar mass of water: "M(H_2O)=18*10^{-3}\frac{kg}{mole}"

Density of water: "\rho=1000\frac{kg}{m^3}"

Volume "V=2.56*10^{[-6}m^3"

Avogadro number: "N_A=6.02*10^{23}mole^{-1}"

So, the number of water molecules:

"N==\frac{10^3*2.56*10^{-6}}{18*10^{-3}}*6.02*10^{23}=0.86*10^{23}"


Problem 2:

a) "\nu=\frac{m}{M}"

Molar mass of ammonia "M=17\frac{g}{mole}"

"\nu=\frac{4.35}{17}=0.26" moles


b) "\nu=\frac{N}{N_A}" whence "N=\nu*N_A"

"N=0.26*6.02*10^{23}=1.57*10^{23}"


c) By the Mendeleev-Clapeiron equation:

"PV=\nu*RT"

P is pressure, "P=10^5Pa"

T is absolute temperature (in degrees of Kelvin), "T=t+273=25+273=298" K

R is universal gas constant "R=8.31\frac{J}{K*mole}" . Then the volume:

"V=\frac{\nu*RT}{P}"

"V=\frac{0.26*8.31*298}{10^5}=6.4" (litres)